Question 4 (10 marks):

Suppose that you have collected data on teenage workers in the fast-food industry, and that you have performed various regression analysis.
Suppose that a regression yields the following results:

Wi =-1+0.3Ai (0.1)

Where W is the hourly wage rate (in dollars) and A is the age (in years). The standard error is in parentheses. A second regression yields this estimate:

Wi = -1 + 0.3 Ai - 0.01(Ai)2

(a) How much is a 20-year-old predicted to earn based on the first regression?

(b) Are you confident that wages actually rise with age? Why?
(c) How much is a 20-year-old predicted to earn based on the second regression?
Answer #1

a) Wi = 1+0.3Ai

at Ai=20 predicted value of Wi =1+0.3(20)

Wi = 7

b) Wi = 1+0.3Ai as coefficient of Ai is 0.3 which is a positive value means that if age increases by 1 year hourly wage would increase by $0.3.

0.3 is also known as slope coefficient

c) Wi = 1+0.3Ai -0.01A^2

at Ai=20 predicted value of Wi =1+0.3(20)-0.01(20)^2

Wi = 1+6-4

Wi=3